# Toward First-Principles Multi-Messenger Predictions: Coupling Nuclear Networks with GR Radiation-MHD in Gmunu

### v1

@@ 3.2 Conserved-to-primitive conversion
Figure 2 shows the averaged error in conserved-to-primitive conversions, sampling $\rho$ and $T$ for selected Lorentz factors $W$, electron fractions $Y_{\rm e}$, and magnetization $\beta_{\rm mag}=P_{\rm gas}/P_{\rm mag}=1000$ (weakly magnetized case). White regions in the lower right reflect temperature recovery issues discussed in Section 3.1. The upper-left region (low density, high temperature) exhibits larger errors for higher Lorentz factors because, in this regime, kinetic energy dominates the conserved variables, making the solution insensitive to thermal energy and causing precision loss due to floating-point cancellation. Tests with stronger magnetic fields (down to $\beta_{\rm mag}=10^{-3}$) show similar behavior and are not shown. Overall, the majority of the parameter space exhibits relative errors below $10^{-8}$, with errors near the edges reaching $\lesssim 10^{-4}$.
@@
3.3 Pure Silicon Burning Test
To verify that the nuclear burning module drives compositions consistently toward NSE, we performed a one-zone calculation starting from pure $\rm{}^{28}Si$. This setup has long been used as a benchmark in nuclear network studies (e.g., Woosley et al. (1973); Hix & Thielemann (1996, 1999); Timmes et al. (2000)), since silicon burning rapidly produces an iron-group dominated composition under high temperatures and densities.
@@ 3.3 Pure Silicon Burning Test
3.4 Newtonian Real Gas Shock Tubes
To verify the stellar EoS implementation, we perform the Newtonian real gas shock tube test introduced by Zingale & Katz (2015), for which exact solutions are available. The simulations are carried out in Cartesian coordinates.
Following test 1 of Zingale & Katz (2015), the initial conditions are
@@ 3.4 Newtonian Real Gas Shock Tubes
Figure 4: Profiles of rest-mass density $\rho$ (upper left), velocity $v$ (upper right), pressure $P$ (lower left), and temperature $T$ (lower right) at $t=8\times 10^{-4}~{\rm s}$ for the Newtonian real gas shock tube test 1 of Zingale & Katz (2015). Red dots: numerical results from Gmunu; black solid lines: exact reference solution. The excellent agreement validates the stellar EoS implementation.
3.5 Newtonian Real Gas Acoustic Pulse in 1D
@@
Following Hasenour & Duffell (2025), we adopt $\rho_{0}=5\times 10^{5}~{\rm g\cdot cm^{-3}}$, $T_{0}=3\times 10^{8}~{\rm K}$, $f_{p}=2$, $\delta_{r}=2\times 10^{7}~{\rm cm}$, and $L_{x}=10^{8}~{\rm cm}$. The fluid is initially composed of pure $\rm{}^{4}He$.
We employ the 2nd-order accurate strong-stability preserving IMEX-SSP2(2,2,2) time integrator (Pareschi & Russo, 2005), the Harten-Lax-van Leer (HLL) Riemann solver, and a 2nd-order monotonized central (MC) limiter (van Leer, 1974). For the nuclear-reacting cases, the 13-species network aprox13 is coupled to the hydrodynamics.
@@
Figure 5: Snapshots of the rest-mass density $\rho$ at six times for the Newtonian real gas acoustic pulse test (Zingale et al., 2019; Hasenour & Duffell, 2025). The upper panel shows pure hydrodynamics, and the lower panel includes nuclear reactions via the 13-species network aprox13. When nuclear burning is included, the expansion dynamics are visibly altered.
@@
As discussed by Hasenour & Duffell (2025), entropy is not conserved when nuclear burning is active; therefore, we monitor convergence using the rest-mass density instead.
@@ 3.6 Detonation front of SNe Ia in 1D
To validate our burning limiter prescription, we reproduce a test problem originally proposed by Papatheodore & Messer (2014), which examines detonations in the outer reactive layers of Type Ia supernovae. This test provides a clean setup for studying the coupling between hydrodynamic shocks and nuclear burning.
@@
The computational domain spans $\left[0,4096\right]~{\rm cm}$ and is discretized with 256 base zones, allowing up to six levels of adaptive mesh refinement (AMR) that track the burning fronts of the $^{12}{\rm C}$ and $^{16}{\rm O}$ mass fractions. At the highest refinement level, the effective grid spacing is $\Delta x=0.5~{\rm cm}$. Reflective boundary conditions are applied at the left boundary (ash side), while outflow boundary conditions are imposed at the right boundary (fuel side). The unperturbed region initially has a uniform rest-mass density and temperature of $\rho=\rho_{0}$ and $T=10^{7}~{\rm K}$, respectively, with equal mass fractions of $^{12}{\rm C}$ and $^{16}{\rm O}$ only. We consider four density setups with $\rho_{0}=5\times 10^{7}$, $1\times 10^{8}$, $5\times 10^{8}$, and $1\times 10^{9}~{\rm g\cdot cm^{-3}}$. To initiate the detonation, we set $T=10^{10}~{\rm K}$ and $v^{x}=10^{9}~{\rm cm\cdot s^{-1}}$ for $x<25~{\rm cm}$.
@@
We perform simulations both with and without the burning limiter mechanism described in Section 2.3.1. In the runs with suppression enabled, nuclear burning is deactivated in zones flagged as shocked, thereby preventing premature or spurious nuclear energy release at the shock front.
@@
Figure 7 shows the rest-mass density profiles at $t=10^{-6}~{\rm s}$ for models with and without burning suppression. Solid lines denote cases where burning is prohibited in shocked regions, while dashed lines correspond to runs where burning is always allowed. Overall, our results qualitatively reproduce the trends reported by Papatheodore & Messer (2014): cases without suppression develop slightly faster-moving shocks due to additional energy deposition from unphysical burning immediately behind the shock. The absolute shock locations differ marginally from those in their study, likely due to differences in the precise output time (their figure does not show results exactly at $t=10^{-6}~{\rm s}$) and in the numerical schemes used.
@@
However, we do not observe the secondary shock structure reported by Papatheodore & Messer (2014). This discrepancy may stem from our use of an IMEX time-integration scheme with direct Newton–Raphson coupling between hydrodynamics and nuclear burning, whereas most previous nuclear hydrodynamics codes evolve the reaction network via operator-split ODE solvers. In addition, differences in numerical dissipation—arising from less diffusive Riemann solvers, reconstruction methods, or the AMR configuration—may also contribute. A more detailed investigation is warranted to assess how different coupling strategies (fully coupled vs. operator-split) influence post-shock burning dynamics and the formation of secondary shocks.
@@
Figure 7: Rest-mass density profiles $\rho$ at $t=10^{-6}~{\rm s}$ for different models with (solid lines) and without (dashed lines) burning suppression at shocks. We qualitatively reproduce the results of Papatheodore & Messer (2014), showing that shocks propagate faster when burning is not suppressed. Minor offsets in shock position are expected due to slight differences in the evaluation time and numerical treatment. Unlike in Papatheodore & Messer (2014), we do not observe the secondary shock feature, which may be related to the IMEX integration scheme used here.
4 Application example: CCSNe in 1D
@@
We consider the collapse of two solar-metallicity progenitors: a 9 M star (Sukhbold et al., 2016) and a 20 M star (Woosley & Heger, 2007). The SFHo EoS (Steiner et al., 2013) is adopted for the NSE regime. Time integration is performed with the IMEX-SSP2(2,2,2) scheme (Pareschi & Russo, 2005), and spatial reconstruction uses the second-order monotonized central (MC) limiter (van Leer, 1974). The computational domain extends over $r\in[0,10^{5}]~{\rm km}$ with $N_{r}=128$ and up to $l_{\max}=13$ mesh-refinement levels, yielding a finest grid spacing of $\Delta r\approx 191~{\rm m}$. Refinement criteria follow Cheong et al. (2023).
@@
Diagnostics are computed as follows. The net neutrino heating rate is estimated by
@@
is the specific enthalpy including only the thermal internal energy. In this expression, $\epsilon$ is the specific internal energy, and $\epsilon_{0}$ is the corresponding zero-point obtained for the same rest-mass density and electron fraction but at zero temperature (Betranhandy & O’Connor, 2020; Cheong et al., 2025). By construction, $\varepsilon_{\mathrm{exp}}>0$ corresponds to the Bernoulli criterion for unbound material. [^1] Strictly speaking, the Bernoulli criterion is not the only possible definition of unbound matter, and can differ from more conservative approaches that account for long-term binding to the proto-neutron star. Nevertheless, it remains the standard diagnostic in CCSN studies and facilitates direct comparison across different works. In the Newtonian limit, equation (32) reduces to the commonly used diagnostic explosion energy in Newtonian CCSN simulations, where the gravitational binding energy is encoded in $u_{t}$. The mass accretion rate at radius $r$ is computed as
@@
[…]
The initialization of composition depends on thermodynamic conditions. In high-density regions where NSE is expected, abundances are computed with the NSE solver (Section 2.3.2). In non-NSE regions, the nuclei present in the progenitor profiles may not match those in the adopted network. We therefore map progenitor abundances to the network species and rescale mass fractions to enforce $\sum X_{l}=1$. The network then relaxes dynamically over the first few timesteps.
@@
All baseline runs without enhanced heating or nuclear burning fail to explode. Our 20 M baseline model is consistent with previous benchmarks obtained using different codes (O’Connor et al., 2018; Ng et al., 2024; Cheong et al., 2024). The 9 M model has also been reported as non-exploding in one-dimensional simulations (Radice et al., 2017). Despite differences in microphysics and gravity treatments—Radice et al. (2017) adopted the stiffer LS220 EoS (Lattimer & Swesty, 1991) and Newtonian gravity with an effective potential (Marek et al., 2006), whereas our models employ full general relativity with the SFHo EoS—our results are qualitatively consistent.
@@ 5 Conclusions
Looking ahead, this work establishes a robust foundation for exploring a wide range of astrophysical explosion scenarios. Future developments will focus on: (i) multidimensional CCSN simulations with self-consistent neutrino heating to capture convection, turbulence, and asphericity; (ii) adaptive nuclear networks that expand dynamically depending on local conditions, eventually incorporating up-to-date libraries such as REACLIB; (iii) systematic studies of nucleosynthesis yields across progenitor models; and (iv) connections to multi-messenger observables, including neutrinos, gravitational waves, and electromagnetic transients. Beyond core-collapse supernovae, the framework is also directly applicable to other explosive phenomena where nuclear burning plays a central role, such as collapsars, white dwarf–neutron star mergers, and white dwarf–white dwarf mergers. Together, these advances will allow Gmunu to become a comprehensive platform for probing the interplay of gravity, neutrino transport, magnetism, and nuclear burning in some of the universe’s most energetic events.
We thank Albino Perego and David Radice for valuable discussions on bridging NSE and non-NSE equations of state. We are also grateful to Gerard Navó and Martin Obergaulinger for their insights regarding the initial setup of the core-collapse supernova tests. We further thank Peter Siu Hei Cheung for helpful general discussions. P.C.-K.C. acknowledges support from the National Science Foundation (NSF) under Grant PHY-2020275, awarded to the Network for Neutrinos, Nuclear Astrophysics, and Symmetries (N3AS). The simulations in this work have been performed on the third UNH supercomputer Marvin, also known as Plasma, which is supported by NSF/Major Research Instrumentatio (MRI) program under grant number AGS-1919310. The simulations have also been performed on the Expanse cluster at San Diego Supercomputer Centre through allocation PHY230104 and PHY230129 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program (Boerner et al., 2023), which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.

### v2

@@ III.2 Conserved-to-primitive conversion
Figure 2 shows the averaged error in conserved-to-primitive conversions, sampling $\rho$ and $T$ for selected Lorentz factors $W$, electron fractions $Y_{\rm e}$, and magnetization $\beta_{\rm mag}=P_{\rm gas}/P_{\rm mag}=1000$ (weakly magnetized case). White regions in the lower right reflect temperature recovery issues discussed in Section III.1. The upper-left region (low density, high temperature) exhibits larger errors for higher Lorentz factors because, in this regime, kinetic energy dominates the conserved variables, making the solution insensitive to thermal energy and causing precision loss due to floating-point cancellation. Tests with stronger magnetic fields (down to $\beta_{\rm mag}=10^{-3}$) show similar behavior and are not shown. Overall, the majority of the parameter space exhibits relative errors below $10^{-8}$, with errors near the edges reaching $\lesssim 10^{-4}$.
@@
III.3 Pure Silicon Burning Test
To verify that the nuclear burning module drives compositions consistently toward NSE, we performed a one-zone calculation starting from pure $\rm{}^{28}Si$. This setup has long been used as a benchmark in nuclear network studies (e.g., Woosley et al. (1973); Hix and Thielemann (1996, 1999); Timmes et al. (2000)), since silicon burning rapidly produces an iron-group dominated composition under high temperatures and densities.
@@ III.3 Pure Silicon Burning Test
The NSE abundances and the asymptotic abundances from the network calculation show slight differences. This arises because the Coulomb physics treated in the NSE solver and in the aprox13 network are not identical. The NSE module incorporates Coulomb corrections following Chabrier and Potekhin (1998). In contrast, aprox13 accounts for Coulomb effects through screening corrections to reaction rates based on Wallace et al. (1982).
@@
III.4 Newtonian Real Gas Shock Tubes
To verify the stellar EoS implementation, we perform the Newtonian real gas shock tube test introduced by Zingale and Katz (2015), for which exact solutions are available. The simulations are carried out in Cartesian coordinates.
Following test 1 of Zingale and Katz (2015), the initial conditions are
@@ III.4 Newtonian Real Gas Shock Tubes
Figure 4: Profiles of rest-mass density $\rho$ (upper left), velocity $v$ (upper right), pressure $P$ (lower left), and temperature $T$ (lower right) at $t=8\times 10^{-4}~{\rm s}$ for the Newtonian real gas shock tube test 1 of Zingale and Katz (2015). Red dots: numerical results from Gmunu; black solid lines: exact reference solution. The excellent agreement validates the stellar EoS implementation.
III.5 Newtonian Real Gas Acoustic Pulse in 1D
@@
Following Hasenour and Duffell (2025), we adopt $\rho_{0}=5\times 10^{5}~{\rm g\cdot cm^{-3}}$, $T_{0}=3\times 10^{8}~{\rm K}$, $f_{p}=2$, $\delta_{r}=2\times 10^{7}~{\rm cm}$, and $L_{x}=10^{8}~{\rm cm}$. The fluid is initially composed of pure $\rm{}^{4}He$.
We employ the 2nd-order accurate strong-stability preserving IMEX-SSP2(2,2,2) time integrator (Pareschi and Russo, 2005), the Harten-Lax-van Leer (HLL) Riemann solver, and a 2nd-order monotonized central (MC) limiter (van Leer, 1974). For the nuclear-reacting cases, the 13-species network aprox13 is coupled to the hydrodynamics.
@@
Figure 5: Snapshots of the rest-mass density $\rho$ at six times for the Newtonian real gas acoustic pulse test (Zingale et al., 2019; Hasenour and Duffell, 2025). The upper panel shows pure hydrodynamics, and the lower panel includes nuclear reactions via the 13-species network aprox13. When nuclear burning is included, the expansion dynamics are visibly altered.
@@
As discussed by Hasenour and Duffell (2025), entropy is not conserved when nuclear burning is active; therefore, we monitor convergence using the rest-mass density instead.
@@ III.6 Detonation front of SNe Ia in 1D
To validate our burning limiter prescription, we reproduce a test problem originally proposed by Papatheodore and Messer (2014), which examines detonations in the outer reactive layers of Type Ia supernovae. This test provides a clean setup for studying the coupling between hydrodynamic shocks and nuclear burning.
@@
The 2nd-order accurate strong-stability preserving IMEX-SSP2(2,2,2) time integrator (Pareschi and Russo, 2005), the Harten-Lax-van Leer (HLL) Riemann solver, and a 2nd-order monotonized central (MC) limiter (van Leer, 1974), are employed in this test. The computational domain spans $\left[0,4096\right]~{\rm cm}$ and is discretized with 256 base zones, allowing up to six levels of adaptive mesh refinement (AMR) that track the burning fronts of the $^{12}{\rm C}$ and $^{16}{\rm O}$ mass fractions. At the highest refinement level, the effective grid spacing is $\Delta x=0.5~{\rm cm}$. Reflective boundary conditions are applied at the left boundary (ash side), while outflow boundary conditions are imposed at the right boundary (fuel side). The unperturbed region initially has a uniform rest-mass density and temperature of $\rho=\rho_{0}$ and $T=10^{7}~{\rm K}$, respectively, with equal mass fractions of $^{12}{\rm C}$ and $^{16}{\rm O}$ only. We consider four density setups with $\rho_{0}=5\times 10^{7}$, $1\times 10^{8}$, $5\times 10^{8}$, and $1\times 10^{9}~{\rm g\cdot cm^{-3}}$. To initiate the detonation, we set $T=10^{10}~{\rm K}$ and $v^{x}=10^{9}~{\rm cm\cdot s^{-1}}$ for $x<25~{\rm cm}$.
@@
We perform simulations both with and without the burning limiter mechanism described in Section II.3.1. In the runs with suppression enabled, nuclear burning is deactivated in zones flagged as shocked, thereby preventing premature or spurious nuclear energy release at the shock front.
@@
Figure 7 shows the rest-mass density profiles at $t=10^{-6}~{\rm s}$ for models with and without burning suppression. Solid lines denote cases where burning is prohibited in shocked regions, while dashed lines correspond to runs where burning is always allowed. Overall, our results qualitatively reproduce the trends reported by Papatheodore and Messer (2014): cases without suppression develop slightly faster-moving shocks due to additional energy deposition from unphysical burning immediately behind the shock. The absolute shock locations differ marginally from those in their study, likely due to differences in the precise output time (their figure does not show results exactly at $t=10^{-6}~{\rm s}$) and in the numerical schemes used.
@@
However, we do not observe the secondary shock structure reported by Papatheodore and Messer (2014). This discrepancy may stem from our use of an IMEX time-integration scheme with direct Newton–Raphson coupling between hydrodynamics and nuclear burning, whereas most previous nuclear hydrodynamics codes evolve the reaction network via operator-split ODE solvers. In addition, differences in numerical dissipation—arising from less diffusive Riemann solvers, reconstruction methods, or the AMR configuration—may also contribute. The piecewise-parabolic method (Colella and Woodward, 1984) used by Papatheodore and Messer (2014) is generally less diffusive and more accurate in smooth regions than the MC reconstruction adopted here. However, because the time-integration scheme and Riemann solvers employed in their work are not documented, a more detailed comparison with Papatheodore and Messer (2014) is not possible here. A more detailed investigation is warranted to assess how different coupling strategies (fully coupled vs. operator-split) influence post-shock burning dynamics and the formation of secondary shocks.
@@
Figure 7: Rest-mass density profiles $\rho$ at $t=10^{-6}~{\rm s}$ for different models with (solid lines) and without (dashed lines) burning suppression at shocks. We qualitatively reproduce the results of Papatheodore and Messer (2014), showing that shocks propagate faster when burning is not suppressed. Minor offsets in shock position are expected due to slight differences in the evaluation time and numerical treatment. Unlike in Papatheodore and Messer (2014), we do not observe the secondary shock feature, which may be related to the IMEX integration scheme used here.
IV Application example: CCSNe in 1D
@@
We consider the collapse of two solar-metallicity progenitors: a 9 M star (Sukhbold et al., 2016) and a 20 M star (Woosley and Heger, 2007). The SFHo EoS (Steiner et al., 2013) is adopted for the NSE regime. Time integration is performed with the IMEX-SSP2(2,2,2) scheme (Pareschi and Russo, 2005), and spatial reconstruction uses the second-order monotonized central (MC) limiter (van Leer, 1974). The computational domain extends over $r\in[0,10^{5}]~{\rm km}$ with $N_{r}=128$ and up to $l_{\max}=13$ mesh-refinement levels, yielding a finest grid spacing of $\Delta r\approx 191~{\rm m}$. Refinement criteria follow Cheong et al. (2023).
@@
Free neutrons and protons are included as advected species but do not participate in the 13-isotope $\alpha$-chain reactions. Their inclusion ensures that the composition can represent an arbitrary $Y_{\rm e}$ when material cooled by neutrino processes enters the low-temperature regime where the network is active. It also allows the NSE solver to operate consistently within the mixed EoS region ($5<T<5.8~{\rm GK}$). A more complete network that explicitly includes neutrons, protons, and heavier nuclei such as $^{56}\mathrm{Fe}$ would remove this limitation; extending the reaction set in this direction is planned for future work.
In the post–bounce phase of a CCSN, the gain region is the layer between the stalled accretion shock and the neutrinosphere where net neutrino heating occurs. In this zone, absorption of electron-type neutrinos dominates over neutrino cooling, resulting in a positive net heating rate. Matter advected through the gain layer can acquire sufficient thermal and pressure support to aid shock revival, making it a key diagnostic for assessing the efficiency of the neutrino-driven explosion mechanism. Following this physical picture, we estimate the net neutrino heating rate by
@@
is the specific enthalpy including only the thermal internal energy. In this expression, $\epsilon$ is the specific internal energy, and $\epsilon_{0}$ is the corresponding zero-point obtained for the same rest-mass density and electron fraction but at zero temperature (Betranhandy and O’Connor, 2020; Cheong et al., 2025). By construction, $\varepsilon_{\mathrm{exp}}>0$ corresponds to the Bernoulli criterion for unbound material. [^1] Strictly speaking, the Bernoulli criterion is not the only possible definition of unbound matter, and can differ from more conservative approaches that account for long-term binding to the proto-neutron star. Nevertheless, it remains the standard diagnostic in CCSN studies and facilitates direct comparison across different works. In the Newtonian limit, equation (32) reduces to the commonly used diagnostic explosion energy in Newtonian CCSN simulations, where the gravitational binding energy is encoded in $u_{t}$.
The mass accretion rate at radius $r$ is computed as
@@
Since spherically symmetric iron-core collapse supernova models usually do not explode without additional treatment, we employ enhanced neutrino heating to trigger shock revival (Blinnikov and Bartunov, 1993; Woosley and Weaver, 1995; Ugliano et al., 2012; Perego et al., 2015; Ertl et al., 2016; Couch et al., 2020; Boccioli et al., 2021; Gogilashvili et al., 2023). Following Gogilashvili et al. (2023), we multiply the energy–momentum and $Y_{\rm e}$ source terms from electron-type neutrinos in the gain region by a constant factor $f_{\rm heat}$. While Gogilashvili et al. (2023) reported explosions of the 20 M model for $f_{\rm heat}\leq 2.4$, here we adopt $f_{\rm heat}=2.8$ for all cases, consistent with Navó et al. (2023). We emphasize that this scheme is artificial and requires calibration; self-consistent explosions require multidimensional simulations. Here it is used solely to probe the role of nuclear burning behind the shock.
The initialization of composition depends on thermodynamic conditions. In high-density regions where NSE is expected, abundances are computed with the NSE solver (Section II.3.2). In non-NSE regions, the nuclei present in the progenitor profiles may not match those in the adopted network. We therefore map progenitor abundances to the network species and rescale mass fractions to enforce $\sum X_{l}=1$. The network then relaxes dynamically over the first few timesteps.
@@
All baseline runs without enhanced heating or nuclear burning fail to explode. Our 20 M baseline model is consistent with previous benchmarks obtained using different codes (O’Connor et al., 2018; Ng et al., 2024; Cheong et al., 2024). The 9 M model has also been reported as non-exploding in one-dimensional simulations (Radice et al., 2017). Despite differences in microphysics and gravity treatments—Radice et al. (2017) adopted the stiffer LS220 EoS (Lattimer and Swesty, 1991) and Newtonian gravity with an effective potential (Marek et al., 2006), whereas our models employ full general relativity with the SFHo EoS—our results are qualitatively consistent.
@@ V Conclusions
Looking ahead, this work establishes a robust foundation for exploring a wide range of astrophysical explosion scenarios. Future developments will focus on: (i) multidimensional CCSN simulations with self-consistent neutrino heating to capture convection, turbulence, and asphericity; (ii) adaptive nuclear networks that expand dynamically depending on local conditions, eventually incorporating up-to-date libraries such as REACLIB and pynucastro (Willcox and Zingale, 2018; Smith et al., 2023; pynucastro development team et al., 2025); (iii) systematic studies of nucleosynthesis yields across progenitor models; and (iv) connections to multi-messenger observables, including neutrinos, gravitational waves, and electromagnetic transients. Beyond core-collapse supernovae, the framework is also directly applicable to other explosive phenomena where nuclear burning plays a central role, such as collapsars, white dwarf–neutron star mergers, and white dwarf–white dwarf mergers. Together, these advances will allow Gmunu to become a comprehensive platform for probing the interplay of gravity, neutrino transport, magnetism, and nuclear burning in some of the universe’s most energetic events.
We thank Albino Perego and David Radice for valuable discussions on bridging NSE and non-NSE equations of state. We are also grateful to Gerard Navó and Martin Obergaulinger for their insights regarding the initial setup of the core-collapse supernova tests. We further thank Peter Siu Hei Cheung for helpful general discussions. P.C.-K.C. acknowledges support from the National Science Foundation (NSF) under Grant PHY-2020275, awarded to the Network for Neutrinos, Nuclear Astrophysics, and Symmetries (N3AS). The simulations in this work have been performed on the third UNH supercomputer Marvin, also known as Plasma, which is supported by NSF/Major Research Instrumentatio (MRI) program under grant number AGS-1919310. The simulations have also been performed on the Expanse cluster at San Diego Supercomputer Centre through allocation PHY230104 and PHY230129 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program (10.1145/3569951.3597559), which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.